# H line shape and radius-luminosity relation in 2.5D FRADO

### v1

@@ 3.2.1 Dependence of line shapes on global parameters
Figures 2, 3, and 4, show the results for the LIL emission line profiles from the entire grid defined in section 2.2. Since the adopted scaling with the incident flux was tested for H line, the results below predominantly apply to this line.
@@
The figure 2 shows the line profiles at representative observer’s viewing angle of 39∘ (Lawrence & Elvis, 2010) for varying black hole masses and for different Eddington ratios as indicated in the legend. For small black hole masses only high values of the Eddington ratios are shown since at too low value of the Eddington ratio even the failed wind is not launched. In this case the lines can form in the disk itself but the inner radius is then not necessarily specified by the FRADO model. We address this issue in the Discussion. As the black hole mass increases, the profiles become broader. This is expected as the gravitational potential well deepens, leading to higher orbital velocities for material forming the BLR, thus reflected in the increasing width of the profiles. Particularly for lower masses the line profiles show a pronounced double-peaked structure. This is indicative of a disk-like dominated geometry of BLR (e.g. Eracleous & Halpern, 1994). This happens for all Eddington rates, when FRADO solution exits. This is not consistent with observational data, as for high Eddington rates and small masses the observed line profiles are single-peaked (e.g. Moran et al., 1996; Laor et al., 1997; Véron-Cetty et al., 2001; Cracco et al., 2016). This suggests that the vertical velocity is underestimated in FRADO for small black hole mass. The feature diminishes for higher masses, where the line profiles are more dominated by a single broad component for high accretion rates, while the double peak is still observed in the low accretion regime. This is consistent with observational trends.
The figure 3 explores the dependence of the H line profile on Eddington ratio while keeping the black hole mass constant in each panel, for a range of masses between 6 to 10, and for the mean viewing angle 39∘. As the accretion rate increases, the H line profiles generally become narrower, with the the Full Width at Half Maximum (FWHM) changing from km/s down to km/s. However, with the rising accretion rate line asymmetry develops, with blueshifted tail. At the vary high accretion rates, close to the Eddington rate, and high masses the line becomes broad again since most of the emission is emitted in an outflow. Lower accretion rates are typically associated with lower radiation pressure and weaker (or nonexistent) outflows, BLR forms closer to the black hole, which means that the gas in the broad-line region (BLR) can be more gravitationally bound and move with higher orbital velocities, leading to broader and double-peaked line profiles. Conversely, at higher accretion rates, the radiation pressure and outflows can dominate, driving winds and possibly increasing turbulence in the BLR. This can introduce more complex velocity structures, and blue asymmetric profiles as radiation pressure pushes gas away, particularly close to the Eddington ratio. This aligns with observational studies where AGNs with high accretion rates (near or above the Eddington limit) tend to have H blueshifted profiles due to radiation pressure-driven winds, while low accretion rate AGNs have symmetric lines due to less interference from such forces. Thus the model is consistent with observational trends (Negrete et al., 2018; Marinello et al., 2020). Strong outflow effects are more clearly observed in the high ionization lines like CIV 1549, which are emitted close to the continuum source and strongly affected by the radiation forces (e.g. Coatman et al., 2017; Martínez-Aldama et al., 2018; Temple et al., 2021). However, those lines forms in the High Ionization Line (HIL) region where line-driving (instead of dust-driving) dominates, and our model does not apply.
@@
The figure 4 displays the line profiles, for the canonic model of AGNs, with the black hole mass set at accreting at Eddington rate (left panel) and for the model corresponding to the mean quasar properties (higher black hole mass and lower Eddington ratio, right panel) from Shen et al. (2011), as a function of the observer’s viewing angle, ranging from 10∘ up to 60∘ (measured from the symmetry axis). The most noticeable effect here is broadening of the line profiles as the viewing angle increases as more of the BLR’s rotational motion is projected along the line of sight. The behavior of the line peak, however, is quite complex and reflects the direction of the outflow. The blueshift mostly dominates but the red wing is very sensitive to the viewing angle.
@@ 3.2.2 Parameters of line shape
Line shapes are most frequently characterized either by FWHM, or by the standard deviation, . These two simple parametrization are not equivalent, and they reflect in particular the predominantly Gaussian or Lorentzian shapes, as discussed in detail by Collin et al. (2006). Since the model provides us with the full line shape, we can study how these shapes are reflected in the usual standard parametrization.
The figure 5 shows the relationship between the standard deviation, , and the of emission lines and other global parameters for a representative fixed viewing angle. As can be seen, there is a positive correlation between these two quantities, suggesting that systems with broader emission lines (higher FWHM) also tend to have larger . Models with higher Eddington ratios generally have smaller and FWHM, while those with lower accretion rates have larger values. It also shows a clear increasing trend of with black hole mass, while the FWHM is more complex, although both parameters show a negative correlation with the Eddington ratio. If we concentrate on the case of a fixed Eddington ratio of 0.3, which is usually considered as a transition from low to high Eddington sources (e.g. Laor & Netzer, 1989; Campitiello et al., 2019), we see that the model for this Eddington ratio gives the of km s-1 for the black hole mass of while for larger mass, is twice that high. It well aligns with the transition between Seyfert 1 galaxies and Narrow Line Seyfert 1 galaxies at FWHM of km s-1 (those are less massive sources) (Peterson, 2011; Paliya et al., 2024), and transition between type A quasars and type B quasars (those are more massive sources) at FWHM of km s-1 (Marziani et al., 2003). The correlation between the FWHM and the accretion rate constrains the results into a defined region with some outliers. The negative relationship between the Eddington ratio and the FWHM (or black hole mass) is more visible in this panel, constraining the low black hole mass objects into the high accretion regime. This result resembles the behavior of the optical Eigenvector 1 (EV1) sequence (Boroson & Green, 1992; Sulentic et al., 2000; Marziani et al., 2003; Shen & Ho, 2014), which correlates the FWHM and optical FeII strength, defined by the equivalent width ratio of FeII emission at 4434-4684 Å over H, =EW(FeII)/EW(H). The linear relationship between and the Eddington ratio (Zamfir et al., 2010), allow us a comparison with the observational findings. However, in the observational case, the region with low accretion rates and low FWHM is populated, which is not observed in our results constrain the predictions into a defined region.
@@
The ratio of the to in line shape studies is known as the shape factor or form factor. This ratio provides information about the shape of the spectral line, and is commonly used to determine whether a line profile is closer to a Gaussian, Lorentzian, or else. For a Gaussian distribution, the shape factor is approximately , and for a Lorentzian distribution, the ratio approaches to zero. For a rectangular function, the ratio is equal 3.46, and a triangular function has a ratio of 2.45. Similarly, for an edge-on rotating ring, it gives a ratio of 2.83 (Collin et al., 2006). The behavior of the shape factor as a function of Eddington ratio, black hole mass, and viewing angle is shown in figure 6.
The results show that as the accretion rate increases, the shape factor decreases, especially for higher mass black holes. At low accretion rates, the shape factor is relatively high, indicating that the line profiles tend to be more Gaussian. This suggests that in these conditions, the broad-line region (BLR) gas has a more uniform velocity distribution, leading to a symmetric, Gaussian-like profile. As the accretion rate increases, the shape factor drops down, especially more pronounced for the higher mass black holes, suggesting that the line profiles become increasingly Lorentzian. This implies that at high accretion rates, the BLR has a more extended velocity distribution, leading to profiles with broader wings, characteristic of Lorentzian shapes. This trend could be attributed to the effect of radiation pressure (and also stronger turbulent motions) at higher accretion rates, leading to a more chaotic pattern of motion in the BLR, which broadens the wings of the emission lines. The change in the gas distribution as a function of the FWHM and Eddington ratio has been widely explored in the EV1 context, where sources with FWHM km s-1 tend to show Lorentzian profiles and high accretion rates, while sources with FWHM km s-1 show Gaussian distributions and low accretion rates (Sulentic et al., 2000). The observational findings agree with our predictions, including the effects of the high radiation pressure (Negrete et al., 2018). The change of the shape factor does not happens at a fixed Eddington ratio but explicitly depends also on the black hole mass. There is, however, a disagreement between the model and the trends known well from the data for small black hole mass . The model clearly underestimates the outflow velocity in low mass systems, and even for high Eddington ratio the line shape is strongly double-peaked, as too small vertical velocity leads to BLR remaining close to the disk surface, with the motion strongly dominated by the Keplerian motion. However, the position of the BLR is consistent with the data trends, and lines are narrow.
@@
Similarly, as the black hole mass increases, for a fixed accretion rate the shape factor tends to decrease. The trend is stronger (starts at smaller black hole masses) for higher accretion rates. For lower black hole masses, the shape factor tends to remain closer to the Gaussian value, indicating a more symmetric velocity distribution of gas around smaller black holes. For more massive black holes, the shape factor drops, especially at high accretion rates. This means that the line profiles become more Lorentzian, likely due to broader wings in the emission lines, caused by more extreme velocity distributions at higher black hole masses. It may seem not intuitive. But for large black hole masses, like typical for quasars, the model predicts that there should be two quasar populations, with either Gaussian or Lorentzian shapes, and the transition is predicted to happen at . This corresponds well to the division of quasars into type A (below thee orange line in Fig 6) and type B (above the orange line in Fig. 6, middle panel). For lower masses the model predicts only Gaussian shape due to too small outflow velocities discussed above.
@@
The shape factor is also affected by the viewing angle, it shows a minimum at intermediate viewing angles (around 20∘ to 25∘), especially for higher black hole masses. At low viewing angles, the shape factor remains relatively close to 2.355, implying that the emission lines appear more Gaussian from a face-on perspective. This suggests that the BLR appears more symmetric when observed from low angles. As the viewing angle increases beyond 30, the shape factor rises again, particularly for more massive black holes. This suggests that for more edge-on views, the profiles again become more Gaussian, possibly due to projection effects that compress the velocity field into a more symmetric profile. The minimum around intermediate angles may arise from a combination of projection effects, and the more complex geometry of the BLR with formation of an outflow structure at higher black hole masses, higher Eddington ratios, where the line-of-sight intersects regions of the BLR with higher velocity dispersion, leading to Lorentzian-like profiles.
@@ 4 Time-delay and R-L relation
4.1 Method of determining the time delay and construction of the R-L relation
From the grid of simulations, we have the 3-D distribution of the clouds representing the geometry of the BLR based on the FRADO model. This distribution allows us to calculate how the radiation emitted in the most central region propagates to the clouds and subsequently reaches the observer. The effect depends not only on the 3-D distribution and emissivity of each cloud but also on the viewing angle of the observer. The time of reprocessing of the radiation within each cloud is neglected since this is a rather fast process in comparison to propagation across the entire BLR zone.
Short pulse from the most central region thus reaches an observer in the form of the extended signal coming from all the clouds which can be stored in the form of the histogram, representing the BLR transfer function. The example of such transfer function is shown in the left panel of figure 7. It assumes that all the clouds are visible to the observer, and no shielding is included. However, we can also experiment with assumptions that some parts of the clouds are not directly visible. With this aim, we divided the clouds into those representing the far side and the near side. In cylindrical coordinates, We define the near side as comprising all clouds within the azimuthal range , which corresponds to the near-facing vertical half of the BLR cake closer to the observer. Conversely, the far side consists of clouds in the range . Likewise, in Cartesian coordinates, taking the observer line-of-sight (arbitrarily due to azimuthal symmetry) in x-z plane, clouds with negative-positive x component are considered as far-near side of the BLR with respect to observer. We thus consider three histograms for each solution.
These histograms were analyzed to extract two key statistical quantities: the average time delay , and the peak time delay . The results are shown in figure 8. The full histograms are always highly asymmetric, in general double-peaked, with sharper first peak from the near side, and more shallow and extended peak from the far side of the BLR. Decomposition shows this clearly.
@@ 4.2 Results
The theoretical R-L relation depends significantly on the options outlined in section 4.1. Strong asymmetry leads to a considerable difference between the peak of the transfer function and the average time delay. Also selective shadowing of the close or far side would change the time delay estimate. We thus study some combinations and compare the results to the observed trends. In observations, the time-delays are usually measured with one of the standard methods, without reconstruction of the transfer function, therefore we do not know a priori which method is more appropriate to compared with the observed R-L.
4.2.1 average time-delays option
We analyze first R-L based on the the average time-delays. The results for a fixed representative viewing angle of 39∘ but all considered values of the black hole mass and Eddington rate are shown in figure 8, left upper corner. The black hole mass is not directly shown in the plot as for a fixed Eddington ratio it can be inferred from the luminosity so that lower luminosity corresponds to lower black hole mass. The results diverge from the observational trends, though the slope and overall distance is promising. Specifically, the average time-delay values displayed a behavior opposite to that expected from the R-L relation regarding the accretion rates. Delays are shorter for low Eddington rate (and they are similar quantitatively to the relation of Bentz et al. (2013), but they are predicted to be longer for higher accretion rates while observations show the opposite (e.g. Du et al., 2015a). This suggests that while the average time delay provides a general sense of the BLR’s overall structure, it may be less sensitive to the actual emission-line response observed in real AGNs, especially when averaged over complex BLR dynamics.
We also explored potential biases that might have arisen due to noise in the observed BLR transfer functions. To address this, we tested removing particles in the simulation with a BLR response ratio below 10% of the peak, which was designed to minimize any spurious contributions from regions of the BLR that are weakly coupled to the continuum variations. Despite this adjustment, the issue with the average time-delays did not resolve.
4.2.2 peak time-delays option

### v3

@@ 3.2.1 Dependence of line shapes on global parameters
Figures 2, 3, and 4 present LIL emission line profiles from the full model grid defined in Section 2.2. The figures are presented in velocity space with positive velocities defined as blueshift. Since the adopted scaling with incident flux was calibrated for the H line, the following results primarily pertain to H.
@@
Figure 2 shows the line profiles at a representative viewing angle of 39∘ (Lawrence & Elvis, 2010) for varying black hole masses and Eddington ratios, as indicated in the legend. For low black hole masses, only high Eddington ratios are shown, since at low accretion rates even the failed wind is not launched. In such cases, lines may form directly in the disk, but the inner BLR radius is not well defined within the FRADO model. This issue is discussed further in Section 6. As the black hole mass increases, the profiles become broader, consistent with deeper gravitational potential wells and higher orbital velocities in the BLR. At lower masses, the line profiles show a pronounced double-peaked structure, indicative of a disk-dominated BLR geometry (e.g., Eracleous & Halpern, 1994). This occurs at all Eddington ratios where a FRADO solution exists. However, this is inconsistent with observations, which show single-peaked profiles for high Eddington ratios and small black hole masses (e.g., Moran et al., 1996; Laor et al., 1997; Véron-Cetty et al., 2001; Cracco et al., 2016). This suggests that vertical velocities may be underestimated by FRADO at low black hole masses. The double-peaked feature diminishes for higher masses, where high accretion rates produce single-peaked, broad profiles, while double peaks remain present in the low accretion regime. This trend is well aligned with observations.
Figure 3 shows the dependence of the H line profile on Eddington ratio, keeping the black hole mass constant in each panel, for values between and and at a representative viewing angle of 39∘. As the accretion rate increases, the H line profiles generally become narrower, with the full width at half maximum (FWHM) decreasing from km/s to km/s. However, the line also develops a blueshifted asymmetry as the accretion rate rises. At very high accretion rates, close to the Eddington limit and for high masses, the line broadens again, as most of the emission originates in an outflow.
At lower accretion rates, radiation pressure is weak or negligible, and outflows are suppressed or absent. The BLR forms closer to the black hole, where gas remains gravitationally bound and moves at higher orbital velocities, producing broader, double-peaked profiles. Conversely, at high accretion rates, radiation pressure and outflows dominate, driving winds and potentially increasing turbulence in the BLR. This leads to more complex velocity fields and blueshifted asymmetric profiles, particularly near the Eddington limit. These results are consistent with observational studies, which show that AGNs with high accretion rates tend to have blueshifted H profiles due to radiation-driven winds, while low accretion rate AGNs exhibit more symmetric lines (Negrete et al., 2018; Marinello et al., 2020).
Strong outflow signatures are more evident in high-ionization lines (HILs) such as C IV , which originate closer to the continuum source and are strongly influenced by radiation forces (e.g., Coatman et al., 2017; Mart´ınez-Aldama et al., 2018; Temple et al., 2021). However, these lines arise in the HIL region where radiation line-driving, rather than dust-driving, dominates; thus, they lie beyond the scope of our model.
@@
Figure 4 shows the H profiles for a canonical AGN model with a black-hole mass of accreting at the Eddington rate (left panel) and for a mean quasar model, with a higher black hole mass and lower Eddington ratio, drawn from Shen et al. (2011) (see the right panel). Profiles are plotted for viewing angles of 10∘ to 60∘, measured from the symmetry axis. The dominant trend is line broadening with increasing inclination, driven by greater projection of the BLR’s rotational velocity along the line of sight. The line core is generally blueshifted, while the red wing remains highly sensitive to viewing angle, reflecting the underlying outflow geometry.
@@ 3.2.2 Parameters of line shape
Line shapes are most frequently characterized by either the FWHM or the standard deviation (). These two standard parametrizations are not equivalent and are typically associated with Gaussian or Lorentzian shapes, as discussed by Collin et al. (2006). Since the model provides the full line shape, we can examine how these shapes map onto these standard parameters.
Figure 5 shows the relationship between the standard deviation () and the FWHM of emission lines, as well as other global parameters, for a representative fixed viewing angle. There is a positive correlation between and FWHM, indicating that broader lines generally exhibit higher values. Higher Eddington ratio models tend to show smaller and FWHM, while lower accretion rates yield broader profiles with larger . A clear increasing trend of with black hole mass is observed, while FWHM shows a more complex dependence; however, we note that both parameters decrease with increasing Eddington ratio.
For a fixed Eddington ratio of 0.3, often considered a transitional value between low and high accretion (e.g., Laor & Netzer, 1989; Campitiello et al., 2019), the model yields km s-1 for a black hole mass of and about twice that for . This aligns well with the transition between Seyfert 1 galaxies and Narrow-Line Seyfert 1 galaxies at FWHM km s-1 (lower mass sources) (Peterson, 2011; Paliya et al., 2024), and the division between type A and type B quasars (more massive) at FWHM km s-1 (Marziani et al., 2003; Panda & Śniegowska, 2024).
The correlation between the FWHM and the accretion rate confines the results to a well-defined region, with just a few outliers. The negative correlation between Eddington ratio and both FWHM and black hole mass is more apparent here, with low-mass black holes constrained to high accretion regimes. This pattern resembles the behavior of the optical Eigenvector 1 (EV1) sequence (Boroson & Green, 1992; Sulentic et al., 2000; Marziani et al., 2003; Shen & Ho, 2014; Panda et al., 2018, 2019), which connects FWHM and optical Fe II strength, quantified by the ratio . The observed linear relationship between and the Eddington ratio (Zamfir et al., 2010) supports a comparison with our model predictions. However, observations do populate the region of low accretion rate and low FWHM, which is not seen in our model, thus confining its predictions to a more limited region.
@@
The ratio of the FWHM to in line profile analysis is commonly referred to as the shape (or form) factor. This ratio provides insight into the shape of the emission line and is often used to assess whether a profile resembles a Gaussian, Lorentzian, or another distribution. For a Gaussian, the shape factor is approximately 2.35; for a Lorentzian, it approaches zero. A rectangular profile yields a value of 3.46, a triangular one gives 2.45, and an edge-on rotating ring yields 2.83 (Collin et al., 2006). Figure 6 shows how the shape factor varies with Eddington ratio, black hole mass, and viewing angle.
The results show that as the accretion rate increases, the shape factor decreases, particularly in higher-mass black holes. At low accretion rates, the shape factor remains relatively high, indicating more symmetric, Gaussian-like profiles. This suggests a more uniform velocity distribution in the BLR under these conditions. As the accretion rate increases, especially in massive systems, the shape factor decreases, indicating increasingly Lorentzian profiles with broader wings. This likely reflects enhanced radiation pressure and turbulence at high accretion rates, leading to more chaotic gas motion in the BLR.
This trend is consistent with EV1 findings, where sources with FWHM ¡ 4000 km s-1 typically show Lorentzian profiles and high accretion rates, while broader lines (FWHM ¿ 4000 km s-1) are associated with Gaussian shapes and lower accretion rates (Sulentic et al., 2000). These observational trends agree with our predictions, including the role of high radiation pressure (Negrete et al., 2018). The transition in shape factor does not occur at a fixed Eddington ratio, but also depends on black hole mass.
However, the model underestimates outflow velocities in low-mass systems (). Even at high accretion rates, the line profiles remain strongly double-peaked due to insufficient vertical motion, keeping the BLR close to the disk and dominated by Keplerian dynamics. Nonetheless, the predicted BLR location remains consistent with observed trends and the lines are appropriately narrow.
@@
Similarly, at a fixed accretion rate, the shape factor decreases with increasing black hole mass. This trend is stronger (and begins at lower black hole masses) for higher accretion rates. For low-mass black holes, the shape factor remains close to the Gaussian value, indicating a more symmetric velocity distribution. As the black hole mass increases, particularly at high accretion rates, the shape factor drops, indicating that line profiles become increasingly Lorentzian. This is likely due to broader wings caused by more extreme velocity fields.
While this may seem counterintuitive, for high black hole masses (e.g., , typical of quasars), the model predicts two distinct quasar populations: one exhibiting Gaussian line shapes and the other Lorentzian. The transition occurs near an accretion rate of , which aligns with the empirical division between type A quasars (below the orange line in Fig. 6) and type B quasars (above the orange line, middle panel). For lower masses, the model predicts only Gaussian shapes, due to the insufficient outflow velocities discussed above.
@@
The shape factor depends on viewing angle, showing a minimum at intermediate inclinations (between and ), especially for higher black hole masses. At low inclinations, it stays near 2.355, indicating a more symmetric BLR appearance and Gaussian-like line profiles when viewed face-on. Beyond , the shape factor increases again, particularly for massive black holes, suggesting that edge-on views yield more Gaussian-like profile. This is possibly due to projection effects compressing the velocity field into a more symmetric form.
The intermediate-angle minimum likely reflects both projection effects and the complex BLR geometry in high-mass, high-Eddington systems. In these cases, the line of sight intersects outflow regions with enhanced velocity dispersion, producing more Lorentzian-like profiles.
@@ 4 Time-delay and R-L relation
4.1 Time delay extraction and R-L construction
From the grid of simulations, we obtained the 3D cloud distribution defining the BLR geometry based on the FRADO model. This distribution allows us to calculate how the central radiation propagates to the clouds and subsequently reaches the observer. The observed response depends on the 3D structure, cloud emissivity, and viewing angle. The reprocessing time within each cloud is neglected, as it is fast compared to light-travel delays across the BLR.
A short pulse from the central source reaches the observer as an extended signal, spread over time by reprocessing in the BLR clouds. This response can be recorded as an emissivity-weighted time-delay histogram, representing the transfer function of the BLR. An example of such a transfer function is shown in the left panel of Figure 7. This case assumes full visibility of all clouds, without any shielding.
To explore obscuration effects, we divided the BLR into near and far sides. In cylindrical coordinates, the near side includes all clouds within the azimuthal range , corresponding to the vertical half of the BLR closest to the observer. The far side covers . In Cartesian coordinates, assuming the observer’s line of sight lies in the x–z plane, clouds with are considered to be on the near side; whereas those with are on the far side. On this basis, we constructed three time-delay histograms for each model: full, near-side only, and far-side only.
These histograms were analyzed to extract two key statistical quantities: the average time delay, , and the peak time delay, . The results are presented in Figure 8. The full histograms are typically asymmetric and often exhibit a double-peaked structure, with a sharper first peak originating from the near side and a broader, shallower second peak from the far side of the BLR. This decomposition makes the contribution from each side clearly visible.
@@ 4.2 Results
The theoretical R–L relation depends significantly on the assumptions described in Section 4.1. Strong asymmetry in the transfer function leads to a noticeable difference between the peak time delay and the average. Additionally, selective visibility of the near or far side can alter the estimated delay. Therefore, we explored various combinations and compare the results to observed trends. In practice, time delays are typically measured using standard methods without reconstructing the full transfer function, so it is not always clear which theoretical definition is most appropriate for comparison with observed R–L relations.
4.2.1 Average time-delay option
We first analyzed the R–L relation based on the average time delays. The results (shown in the upper-left panel of Figure 8) correspond to a fixed representative viewing angle of , covering all considered black hole masses and Eddington ratios. The black hole mass is not shown explicitly, but for a fixed Eddington ratio, it can be inferred from the luminosity, whereby lower luminosities correspond to lower black hole masses.
The results deviate from observational trends, although the slope and overall location relative to the Bentz et al. (2013) relation are promising. Specifically, the average time delays exhibit behavior opposite to that seen in observations with respect to accretion rate. The model predicts shorter delays for low Eddington ratio sources, consistent with Bentz et al. (2013), but longer delays for high accretion rates, in contrast to observational findings (e.g., Du et al., 2015). This suggests that while the average delay reflects the overall BLR structure, it may not accurately capture the emission-line response seen in reverberation mapping, particularly under complex BLR dynamics.
We also explored potential biases that could arise from noise in observed BLR transfer functions. To test this, we removed simulated particles with response strengths below 10% of the peak, aiming to exclude weakly responding regions that may contribute spurious delays. However, this adjustment did not resolve the discrepancy; specifically, the model’s prediction of increasing average time delays with higher accretion rates, which is not observed in reverberation data. This discrepancy suggests that the issue is likely intrinsically linked to the way the average delay reflects the BLR structure in the model.
4.2.2 Peak time-delay option